Please help me with these:
Sole the equations
1. 5-3(2b-3)=-4
2. 6(4x+2)=56+2x
Thanks.

Found 2 solutions by TimothyLamb, stanbon:
Answer by TimothyLamb(4379)

One question per post please:
---
5 - 3(2b - 3)= -4
5 - 6b + 9 = -4
6b = 5 + 9 + 4
b = (5 + 9 + 4)/6
b = 3
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php

Answer by stanbon(75887)

Solve the equations
1. 5-3(2b-3)=-4
5-6b+9 = -4
-6b + 14 = -4
6b = 14 + 4
6b = 18
b = 3
-------------
2. 6(4x+2)=56+2x
24x + 12 = 56 + 2x
---
22x = 44
x = 2
--------
Cheers,
Stan H.
--------